# Unmasking the Shadows: Pinpoint the Implementations of Anti-Dynamic Analysis Techniques in Malware Using LLM

### v1

@@ III-C Problem Statement
In this paper, we decide to perform our analysis at the basic block ($B$) level, and correspondingly, the concept of ”near” will be a small number of $B$ away from the $B$ where the breakpoint is located, on the control flow graph (CFG). Specifically, we want to place the breakpoints at the $B$s that are parts of the TADA implementations.
@@ IV-A Overview
What is the granularity of the analysis? The breakpoints are usually placed at the beginning of instructions so that the granularity of the breakpoints is at the instruction level. However, for the sake of detecting TADA implementations, instruction level could be improper, as the amount of information carried in an single instruction could be too little. The most obvious example are API calls, which usually involves more than one instructions if any arguments are presented. On the other hand, function-level could be too large, as placing a breakpoint at a function could be futile, in worse cases requiring human analyst manually inspecting thousands of instructions. Therefore, to include sufficient information for detection while keeping the definition of the breakpoint location reasonably precise, our detection granularity is at the basic block ($B$) level.
What input fits LLM well? This paper focuses on reverse engineering of binary programs, which is clearly inappropriate to be directly integrated into the input prompt. Essentially, both open-source and proprietary LLMs are trained using natural languages in addition to some program languages so that the LLMs cannot handle reverse engineering tasks from scratch. Common strategies to address this type of challenge include few-shot learning and fine-tuning, but given the diversity of the TADA implementations, such strategies face the challenge of collecting high-quality training datasets. Therefore, to ensure the LLM is well exploited, instead of providing hex strings of raw binary or assembly code, we first create necessary features to find TADA implementations for each $B$, that are described in the form of natural language. Despite our $B$ level granularity, it does not mean that our analysis is at $B$ level only, while we leverage control flow and data flow analysis to extract the features we need.
Workflow summary. At a higher level, we first conduct program analysis to extract the features that are necessary to determine whether a $B$ is TADA-related or not. The necessity and the selection of the features are based on domain knowledge, which will be explained in detail in the following subsections. Those features will be then used to build prompts and subsequently sent to LLM for the final output. Although the output of our workflow is a binary decision for each $B$, we prompt the LLM so that a response of ratings (of how related the $B$ is to TADA implementations) is generated for each $B$.
@@
The major steps are illustrated in Figure 2. Starting from a malware sample, we first construct the control flow graph (CFG) and extract all the $B$s from the executable. $B$s consist of instructions, from which along with the static analysis over the whole binary we extract three kinds of features: assembly features, API call features, and string features. After the features are extracted, they will be used to craft the prompt, which will be then sent to the LLM, querying whether the basic block is part of the implementation of TADA and suitable for a breakpoint.
IV-B Sample unpacking and $B$ extraction
@@
To this point, if the original IAT can be parsed, constructing CFG and dumping the $B$s are fairly straightforward. Existing tools include but are not limited to IDA, Angr, etc.
@@ IV-C Feature Construction: Assembly Feature
Assembly features are features extracted solely from the assembly language, which can be acquired directly from the extracted $B$s. In particular, we are looking for two specific pieces of information: mnemonics of instructions and memory accesses.
In assembly language, mnemonics of instruction are the most semantic-rich component, and the presence of certain mnemonics can be a strong indicator of TADAs. For example, if a $B$ contains both pushf and popf instructions, then it is extremely suspicious to detect a debugger. For another example, if a $B$ contains hardware querying instructions such as cpuid, then it could be detected if a VM exists. As for memory accesses, we particularly look for memory accesses through segment registers. On X86 Windows host, many processes-related information can be accessed through segment registers. For example, the process environment block (PEB) can be accessed through fs register, which contains various information that can indicate whether a process is being debugged or traced.
Accordingly, for each $B$, we iterate through each instruction and check whether the instruction has uncommon, TADA-related mnemonics. In our experiments, we include in total of 15 uncommon mnemonics, which are shown in Appendix A. If any uncommon mnemonics are found, we will report this finding in the prompt, such as:
@@ IV-D Feature Construction: String Feature
One key motivation for using LLM is that the string can carry many clues regarding whether a $B$ is TADA-related. For example, as shown in Figure 3, if a $B$ refers to a string such as VirtualBox, then obviously the $B$ is very suspicious of involving TADA. Accordingly, we include string features, such as:
@@
Figure 4 shows the overall logic of our string deobfuscation process. First, we check whether the data reference is a plain string: a string saved in plain text format with no obfuscation. If not, we will emulate the function whenever we suspect the data reference is referring to data that could be a string. We have two major criteria to determine whether the emulation will be performed: 1) whether a loop, especially a single $B$ loop is presented; and 2) whether there are consecutive mov instructions (more than 5) that are moving data from data sections to the stack. The intuitions behind the two criteria are simple: for encoded strings, a loop is extremely likely to be presented, if not inevitable, as decoding data almost always requires loops; for stack and stack encoded strings, consecutive mov instructions must present to set up the stack. If any of the criteria is met, we will then do the emulation to find the strings. After a string is found, it will first be used to construct a string feature as shown above, and then be recorded for later use.
@@ IV-E Feature Construction: API Call Feature
As the name suggested, the API call feature of a $B$ reflects all the API calls in the $B$. Needless to mention, APIs called during malware execution are one of the most important information to understand the behavior of the sample. For our purposes, API calls are indeed critical as well, as numerous TADAs can be implemented through APIs. For example, the most famous single API that can be used to do anti-debugger is IsDebuggerPresent, which will return the value of the BeingDebugged flag in PEB  [^1] https://learn.microsoft.com/en-us/windows/win32/api/winternl/ns-winternl-peb. For another example, in many anti-sandbox techniques, malware authors will conduct detection based on computer and username, which can be very difficult to acquire without using the Windows APIs such as GetComputerName. Therefore, APIs is extremely important for detecting TADA.
@@
In our API call feature, we not only include the API name, but also the arguments passed to the API. Shown in Figure 5 is an example of API call in a $B$ calling MessageBoxA. Since the call instruction refers DS register, after checking the IAT, we will be able to find out that this call is calling MessageBoxA. Subsequently, based on the documentation of the MessageBoxA, we know that the 4 pushed data are the 4 arguments of the call.
@@
Indirect calls are extremely challenging, since our feature extraction only uses static analysis. However, there are still simple cases where calls may be resolved, as long as the register value at the call site can be inferred. To do so, if we find an indirect call instruction in a $B$, we will construct the data flow graph (DFG) for the function containing the $B$. Then we trace the data flow backward from the indirect call site to see if a live variable with a concrete value can be found (e.g. value from static data sections). If so, we will check if the value is a valid address for an API, in a similar way as described earlier for direct calls.
@@ IV-F Leveraging LLM
Output. At first glance, it may seem that the most intuitive model of this problem is binary classification: having the LLMs to output a binary verdict of whether a $B$ is TADA related or not. However, as described in Section III-B, identifying implementation of TADA is a very challenging task and is very prone to false positive, so that the best scenario for a human analyst is to have quantified results which can give them a sense of which $B$ should be prioritized for investigation. Instead of prompting the LLM to output a binary result, we prompt for a rating. The LLM will essentially output a number from 0 to 10, reflecting how related a $B$ may be related to the implementation of TADA.
@@
We set a threshold at 7 for positive results, that is, $B$s that receive a rating of 7 are considered to be related to TADA implementation.
@@ V Evaluation
Our implementation is based on the workflow introduced in Section IV, which will be elaborated in the next subsection. The output of our workflow is binary for each of the $B$: whether a breakpoint should be placed at the beginning of the $B$. Every $B$ that receive an LLM rating of 7 or above is considered as positive $B$ and a breakpoint shall be placed.
@@ V-B Can our method identify well-known TADAs?
Despite the diversity of TADA mentioned in Section III-B, there still exists many popular and well-known TADAs, which are commonly adopted by malware authors. Thus, we first need to ensure that our method can identify the $B$s implementing these popular TADAs before we explore those interesting tricky ones. We built a set of programs, each of which includes an implementation of one well-known TADA. Instead of implementing everything from scratch, we adopted publicly available implementations of these TADAs from a popular GitHub repository of malware anti-analysis techniques, al-khaser.
@@
Since all the TADAs are self-implemented, the compiled programs are not packed. All 164 programs are analyzed using the workflow described in Section IV, except for the malware unpacking components. We consider an implementation of TADA to be detected if and only if at least one $B$ that is a part of the implementation of the TADA is reported by our tool as positive (e.g., LLM rating greater than 7). Therefore, in this experiment, the maximum number of positive $B$s can be detected are far more than 164, in cases where the implementation of a TADA involves many $B$s. However, for the sake of answering the research question, we will discuss our result in this experiment in terms of the number of the TADA implementation detected, not the number of $B$s.
@@
In summary, out of all the 164 TADA implementations, we have successfully detected at least one $B$ in 144 of them, yielding a detection rate of 87.80%. To conduct the root causes analysis for the false negatives, we start with investigating the detection rate per category.
Shown in Figure 6 is the detection rates per-tatic category. It is noteworthy that out of 35 debugger evasion TADAs, only 24 of them are detected and 11 of them are false negatives, rendering a detection rate of merely 68.57%. This number is particularly low compared to the average detection rate mentioned earlier. After we studied the false negatives, we found one of the major reasons is that the indirect calls are not accurately resolved by our tool, which can be clearly reflected in Figure 7. The second reason is that some API calls are not directly related to the anti-dynamic-analysis, causing the LLM to give conservative ratings for those $B$s calling those APIs. Back to the causes of the particularly low detection rate of the debugger evasion category, the reason to be blamed is that most of debugger evasion TADAs can be implemented purely using single or few APIs.
@@ V-C1 Anti-emulation in Raspberry Robin
We have done analysis on samples of Raspberry Robin ($R$), which is a very active malware family. In late 2023, analysts found some $R$ samples were adopting an interesting method to detect Windows Defender Emulation engines [24].
Running our tool against the sample with anti-emulation techniques, we identify a $B$ that is not so obvious to be related to TADA. This $B$ mainly consists of the following features:
@@ V-C2 Disk Serial Number Check in Zebrocy
Running our tool against the sample shown in the blog, we are able to detect the first $B$ of the implementation of this TADA technique. The $B$ mainly contains following features:
@@
We confirmed our findings based on the blog post. It is shown that this is the $B$ querying the disk serial number, followed by a series of comparisons with known disk serial numbers used by sandboxes.
@@ V-C3 Comprehensive Hardware Checks in Trickbot
Our tool study the sample with hash listed in [19], and found several $B$s that are directly related to TADA. For examples, we found a $B$ contains code querying WMI:
@@
In addition, this sample also tried to detect virtual environments. One example is shown in the following $B$:
@@ V-C4 Analysis Tool Searching in Mustung Panda
During our real-world sample analysis, we have also seen some old-school yet still very effective methods against dynamic analysis. One example is Mustung Panda family, which adopted multiple TADA. Firstly, Mustung Panda use CPUID instruction to query processor information. The $B$ which performs this technique will have features as follows:
@@
Besides, Mustung panda samples not only check for hardware such as processor, but also search for reverse engineering tools. One $B$ implementing such technique has features as follows:
@@
Both $B$ are among the detected $B$s by our tool with high confidence.
@@ V-D How much time our method may save for the human analysts?
This is a difficult question to answer directly in quantitative manner. Therefore, instead we evaluate this question indirectly by investigating the number of $B$s that our tool can find out of a program.
@@
We first dissemble the samples and compute the total number of $B$s of each sample. On average, the total number of $B$s in a sample is 6869.31, and after analyzing all 30 samples using our tool, we found that on average we can find 12.57 $B$s positive for TADA implementations. In these 30 samples, the number of detected $B$s is significantly smaller than that of the total number of $B$s, whose ratio is about 0.17%. We believe this ratio indicates that finding code of TADA is really finding a needle in a haystack, and our tool should be able to save a good amount of time for the human analysts.
@@ VI-B Limitations
Limited LLM prompt engineering. In this work, we primarily focus on specifying the definition of the TADA in the prompt, and we did not adopt any advanced prompt engineering methods. One future work is to use dynamic few-shot learning during the prompt crafting phase, so that an example set of features of TADA related $B$ can be added in the prompt. This will enable the LLM to learn from the ground truth of existing known TADA $B$s.

### v2

@@ III-C Problem Statement
In this paper, we decide to perform our analysis at the basic block (BB) level, and correspondingly, the concept of ”near” will be a small number of BB away from the BB where the breakpoint is located, on the control flow graph (CFG). Specifically, we want to place the breakpoints at the BBs that are parts of the TADA implementations.
@@ IV-A Overview
What is the granularity of the analysis? The breakpoints are usually placed at the beginning of instructions so that the granularity of the breakpoints is at the instruction level. However, for the sake of detecting TADA implementations, instruction level could be improper, as the amount of information carried in an single instruction could be too little. The most obvious example are API calls, which usually involves more than one instructions if any arguments are presented. On the other hand, function-level could be too large, as placing a breakpoint at a function could be futile, in worse cases requiring human analyst manually inspecting thousands of instructions. Therefore, to include sufficient information for detection while keeping the definition of the breakpoint location reasonably precise, our detection granularity is at the basic block (BB) level.
What input fits LLM well? This paper focuses on reverse engineering of binary programs, which is clearly inappropriate to be directly integrated into the input prompt. Essentially, both open-source and proprietary LLMs are trained using natural languages in addition to some program languages so that the LLMs cannot handle reverse engineering tasks from scratch. Common strategies to address this type of challenge include few-shot learning and fine-tuning, but given the diversity of the TADA implementations, such strategies face the challenge of collecting high-quality training datasets. Therefore, to ensure the LLM is well exploited, instead of providing hex strings of raw binary or assembly code, we first create necessary features to find TADA implementations for each BB, that are described in the form of natural language. Despite our BB level granularity, it does not mean that our analysis is at BB level only, while we leverage control flow and data flow analysis to extract the features we need.
Workflow summary. At a higher level, we first conduct program analysis to extract the features that are necessary to determine whether a BB is TADA-related or not. The necessity and the selection of the features are based on domain knowledge, which will be explained in detail in the following subsections. Those features will be then used to build prompts and subsequently sent to LLM for the final output. Although the output of our workflow is a binary decision for each BB, we prompt the LLM so that a response of ratings (of how related the BB is to TADA implementations) is generated for each BB.
@@
The major steps are illustrated in Figure 2. Starting from a malware sample, we first construct the control flow graph (CFG) and extract all the BBs from the executable. BBs consist of instructions, from which along with the static analysis over the whole binary we extract three kinds of features: assembly features, API call features, and string features. After the features are extracted, they will be used to craft the prompt, which will be then sent to the LLM, querying whether the basic block is part of the implementation of TADA and suitable for a breakpoint.
IV-B Sample unpacking and BB extraction
@@
To this point, if the original IAT can be parsed, constructing CFG and dumping the BBs are fairly straightforward. Existing tools include but are not limited to IDA, Angr, etc.
@@ IV-C Feature Construction: Assembly Feature
Assembly features are features extracted solely from the assembly language, which can be acquired directly from the extracted BBs. In particular, we are looking for two specific pieces of information: mnemonics of instructions and memory accesses.
In assembly language, mnemonics of instruction are the most semantic-rich component, and the presence of certain mnemonics can be a strong indicator of TADAs. For example, if a BB contains both pushf and popf instructions, then it is extremely suspicious to detect a debugger. For another example, if a BB contains hardware querying instructions such as cpuid, then it could be detected if a VM exists. As for memory accesses, we particularly look for memory accesses through segment registers. On X86 Windows host, many processes-related information can be accessed through segment registers. For example, the process environment block (PEB) can be accessed through fs register, which contains various information that can indicate whether a process is being debugged or traced.
Accordingly, for each BB, we iterate through each instruction and check whether the instruction has uncommon, TADA-related mnemonics. In our experiments, we include in total of 15 uncommon mnemonics, which are shown in Appendix A. If any uncommon mnemonics are found, we will report this finding in the prompt, such as:
@@ IV-D Feature Construction: String Feature
One key motivation for using LLM is that the string can carry many clues regarding whether a BB is TADA-related. For example, as shown in Figure 3, if a BB refers to a string such as VirtualBox, then obviously the BB is very suspicious of involving TADA. Accordingly, we include string features, such as:
@@
Figure 4 shows the overall logic of our string deobfuscation process. First, we check whether the data reference is a plain string: a string saved in plain text format with no obfuscation. If not, we will emulate the function whenever we suspect the data reference is referring to data that could be a string. We have two major criteria to determine whether the emulation will be performed: 1) whether a loop, especially a single BB loop is presented; and 2) whether there are consecutive mov instructions (more than 5) that are moving data from data sections to the stack. The intuitions behind the two criteria are simple: for encoded strings, a loop is extremely likely to be presented, if not inevitable, as decoding data almost always requires loops; for stack and stack encoded strings, consecutive mov instructions must present to set up the stack. If any of the criteria is met, we will then do the emulation to find the strings. After a string is found, it will first be used to construct a string feature as shown above, and then be recorded for later use.
@@ IV-E Feature Construction: API Call Feature
As the name suggested, the API call feature of a BB reflects all the API calls in the BB. Needless to mention, APIs called during malware execution are one of the most important information to understand the behavior of the sample. For our purposes, API calls are indeed critical as well, as numerous TADAs can be implemented through APIs. For example, the most famous single API that can be used to do anti-debugger is IsDebuggerPresent, which will return the value of the BeingDebugged flag in PEB  [^1] https://learn.microsoft.com/en-us/windows/win32/api/winternl/ns-winternl-peb. For another example, in many anti-sandbox techniques, malware authors will conduct detection based on computer and username, which can be very difficult to acquire without using the Windows APIs such as GetComputerName. Therefore, APIs is extremely important for detecting TADA.
@@
In our API call feature, we not only include the API name, but also the arguments passed to the API. Shown in Figure 5 is an example of API call in a BB calling MessageBoxA. Since the call instruction refers DS register, after checking the IAT, we will be able to find out that this call is calling MessageBoxA. Subsequently, based on the documentation of the MessageBoxA, we know that the 4 pushed data are the 4 arguments of the call.
@@
Indirect calls are extremely challenging, since our feature extraction only uses static analysis. However, there are still simple cases where calls may be resolved, as long as the register value at the call site can be inferred. To do so, if we find an indirect call instruction in a BB, we will construct the data flow graph (DFG) for the function containing the BB. Then we trace the data flow backward from the indirect call site to see if a live variable with a concrete value can be found (e.g. value from static data sections). If so, we will check if the value is a valid address for an API, in a similar way as described earlier for direct calls.
@@ IV-F Leveraging LLM
Output. At first glance, it may seem that the most intuitive model of this problem is binary classification: having the LLMs to output a binary verdict of whether a BB is TADA related or not. However, as described in Section III-B, identifying implementation of TADA is a very challenging task and is very prone to false positive, so that the best scenario for a human analyst is to have quantified results which can give them a sense of which BB should be prioritized for investigation. Instead of prompting the LLM to output a binary result, we prompt for a rating. The LLM will essentially output a number from 0 to 10, reflecting how related a BB may be related to the implementation of TADA.
@@
We set a threshold at 7 for positive results, that is, BBs that receive a rating of 7 are considered to be related to TADA implementation.
@@ V Evaluation
Our implementation is based on the workflow introduced in Section IV, which will be elaborated in the next subsection. The output of our workflow is binary for each of the BB: whether a breakpoint should be placed at the beginning of the BB. Every BB that receive an LLM rating of 7 or above is considered as positive BB and a breakpoint shall be placed.
@@ V-B Can our method identify well-known TADAs?
Despite the diversity of TADA mentioned in Section III-B, there still exists many popular and well-known TADAs, which are commonly adopted by malware authors. Thus, we first need to ensure that our method can identify the BBs implementing these popular TADAs before we explore those interesting tricky ones. We built a set of programs, each of which includes an implementation of one well-known TADA. Instead of implementing everything from scratch, we adopted publicly available implementations of these TADAs from a popular GitHub repository of malware anti-analysis techniques, al-khaser.
@@
Since all the TADAs are self-implemented, the compiled programs are not packed. All 164 programs are analyzed using the workflow described in Section IV, except for the malware unpacking components. We consider an implementation of TADA to be detected if and only if at least one BB that is a part of the implementation of the TADA is reported by our tool as positive (e.g., LLM rating greater than 7). Therefore, in this experiment, the maximum number of positive BBs can be detected are far more than 164, in cases where the implementation of a TADA involves many BBs. However, for the sake of answering the research question, we will discuss our result in this experiment in terms of the number of the TADA implementation detected, not the number of BBs.
@@
In summary, out of all the 164 TADA implementations, we have successfully detected at least one BB in 144 of them, yielding a detection rate of 87.80%. To conduct the root causes analysis for the false negatives, we start with investigating the detection rate per category.
Shown in Figure 6 is the detection rates per-tatic category. It is noteworthy that out of 35 debugger evasion TADAs, only 24 of them are detected and 11 of them are false negatives, rendering a detection rate of merely 68.57%. This number is particularly low compared to the average detection rate mentioned earlier. After we studied the false negatives, we found one of the major reasons is that the indirect calls are not accurately resolved by our tool, which can be clearly reflected in Figure 7. The second reason is that some API calls are not directly related to the anti-dynamic-analysis, causing the LLM to give conservative ratings for those BBs calling those APIs. Back to the causes of the particularly low detection rate of the debugger evasion category, the reason to be blamed is that most of debugger evasion TADAs can be implemented purely using single or few APIs.
@@ V-C1 Anti-emulation in Raspberry Robin
We have done analysis on samples of Raspberry Robin (RR), which is a very active malware family. In late 2023, analysts found some RR samples were adopting an interesting method to detect Windows Defender Emulation engines [24].
Running our tool against the sample with anti-emulation techniques, we identify a BB that is not so obvious to be related to TADA. This BB mainly consists of the following features:
@@ V-C2 Disk Serial Number Check in Zebrocy
Running our tool against the sample shown in the blog, we are able to detect the first BB of the implementation of this TADA technique. The BB mainly contains following features:
@@
We confirmed our findings based on the blog post. It is shown that this is the BB querying the disk serial number, followed by a series of comparisons with known disk serial numbers used by sandboxes.
@@ V-C3 Comprehensive Hardware Checks in Trickbot
Our tool study the sample with hash listed in [19], and found several BBs that are directly related to TADA. For examples, we found a BB contains code querying WMI:
@@
In addition, this sample also tried to detect virtual environments. One example is shown in the following BB:
@@ V-C4 Analysis Tool Searching in Mustung Panda
During our real-world sample analysis, we have also seen some old-school yet still very effective methods against dynamic analysis. One example is Mustung Panda family, which adopted multiple TADA. Firstly, Mustung Panda use CPUID instruction to query processor information. The BB which performs this technique will have features as follows:
@@
Besides, Mustung panda samples not only check for hardware such as processor, but also search for reverse engineering tools. One BB implementing such technique has features as follows:
@@
Both BB are among the detected BBs by our tool with high confidence.
@@ V-D How much time our method may save for the human analysts?
This is a difficult question to answer directly in quantitative manner. Therefore, instead we evaluate this question indirectly by investigating the number of BBs that our tool can find out of a program.
@@
We first dissemble the samples and compute the total number of BBs of each sample. On average, the total number of BBs in a sample is 6869.31, and after analyzing all 30 samples using our tool, we found that on average we can find 12.57 BBs positive for TADA implementations. In these 30 samples, the number of detected BBs is significantly smaller than that of the total number of BBs, whose ratio is about 0.17%. We believe this ratio indicates that finding code of TADA is really finding a needle in a haystack, and our tool should be able to save a good amount of time for the human analysts.
@@ VI-B Limitations
Limited LLM prompt engineering. In this work, we primarily focus on specifying the definition of the TADA in the prompt, and we did not adopt any advanced prompt engineering methods. One future work is to use dynamic few-shot learning during the prompt crafting phase, so that an example set of features of TADA related BB can be added in the prompt. This will enable the LLM to learn from the ground truth of existing known TADA BBs.